OpenStudy (anonymous):

Solve the following system of equations. 3y = 9 9y + 2x = -12

OpenStudy (anonymous):

A. (1/3, 15/2) B. (-1/3, - 15/2) C. (8, 8) D. (-39/2, 3)

terenzreignz (terenzreignz):

Well, let's look at the simpler things in life :) 3y = 9, how would you simplify that?

OpenStudy (anonymous):

Divid by 3 on both sides?

terenzreignz (terenzreignz):

That's right, and what does that give you?

OpenStudy (anonymous):

3

terenzreignz (terenzreignz):

3? that is not an equation ;)

terenzreignz (terenzreignz):

it gives you y = 3, right?

OpenStudy (anonymous):

yeah d;

terenzreignz (terenzreignz):

well, now that you have a y-value, you can simply substitute it on the second equation giving you a single equation in one variable go ahead :)

OpenStudy (anonymous):

3(3y) + 2x = -12 3(9) + 2x = -12 x = -39/2... So, 3y = 9 y = 3 D is the answer..

terenzreignz (terenzreignz):

or you could do that ^

OpenStudy (anonymous):

Thanks you guys ^.<

OpenStudy (anonymous):

Welcome dear..